# How to Evaluate Monocular Depth Estimation?

### v1

@@ 5 Sensitivity Aligned Composition
Our sensitivity analysis of existing metrics reveals that no existing metrics align well with human judgment. How do we create a metric that does align well? A well-aligned metric can be useful for human-facing generative applications.
@@
Similar to WKDR (Zoran et al. (2015)) and boundary F1 score (Bochkovskii et al. (2025)), we consider the relation between patches (sets of pixels) within the depth map. Given patches $p$ and $q$ we calculate the normal of the surface described by each patch and the angle between these normals $\angle(\hat{n}_{p},\hat{n}_{q}),\angle(n_{p},n_{q})$ for both ground truth and prediction. We take the average difference over many such patches,
@@
Our selection of patches is motivated by the following procedure: first uniformly select a pixel $I$ within the image, then select a pixel $J$ uniformly within a neighborhood of $I$, finally compute the error. To approximate this expectation over a multivariate uniform distribution, we use the first $m$ elements of the Sobol sequence. This provides a deterministic algorithm for computing the metric. In the appendix we show that 1 million samples is sufficient when computing this metric on standard RGB-D datasets. We also take the average over multiple scales by downsampling the image by a factor $k$ and computing the relative normals. The following results use a neighborhood radius of 32 and scale factors $[1,2,4,8]$. We compute the normals using the cross product of vectors from the top/bottom and left/right pixels adjacent to a central pixel. For datasets with noisy ground truth, it may be preferable to compute normals with a plane of best fit.
@@
Examples of SAWA-H on Real Predictions The SAWA-H (with RelNormal) metric provides different rankings among depth predictions than the AbsRel metric. In figure 7 we visualize this difference for the predictions of Depth Pro and Metric3DV2 on an Infinigen scene using error maps. This is mostly straightforward, though to compute the error map for the boundary f1 score we divide the error equally among all edge pixels where the ground truth and predicted depth map disagree. Though Metric3D achieves a smaller AbsRel error, its SAWA-H error is substantially larger. This is due to the fact that the Depth Pro model made a mistake judging the scale of the iceberg and orientation of the camera relative to ocean. Hence, the iceberg and ocean cannot both be properly aligned, causing high error. However, as seen in the projected contour plot, Depth Pro better approximates the local geometry and has a lower SAWA-H error score. We present more examples with predictions on Infinigen and iBims (Koch et al., 2019) in Appendix  E.
Performance of MDE Methods under SAWA-H We use SAWA-H (with RelNormal) to evaluate 10 state-of-the-art monocular depth estimation methods: DepthAnything(Yang et al., 2024a), DepthAnythingV2(Yang et al., 2024b), GeoWizard(Fu et al., 2024), Marigold(Ke et al., 2024), MiDaS(Birkl et al., 2023), DepthPro(Bochkovskii et al., 2025), Metric3DV2(Hu et al., 2024), MoGe-2(Wang et al., 2025b), UniDepthV2(Piccinelli et al., 2025b), and ZoeDepth(Bhat et al., 2023). Evaluation is conducted on 9 commonly used benchmarks: DDAD(Guizilini et al., 2020), DIODE(Vasiljevic et al., 2019), ETH3D(Schops et al., 2019), HAMMER(Jung et al., 2023), iBims-1(Koch et al., 2019), KITTI(Uhrig et al., 2017), NYUv2(Silberman et al., 2012), Sintel(Butler et al., 2012) and Spring(Mehl et al., 2023).
@@
Fig. 8 presents metric values of each method averaged over 9 datasets. Performance of methods on each dataset are reported in Appendix G. Note that Boundary F1 is one component of SAWA-H but some datasets do not provide valid depth value on pixels near boundaries, preventing SAWA-H from properly penalizing blurry boundaries on those datasets. We additionally report performance of methods averaged on datasets with accurate depth across boundaries in Appendix G.
@@ 7 Acknowledgments
Appendix A Perturbation Algorithm Of Controllable Intensity
@@
Appendix B More on Collecting Data of Human Judgment
@@ Appendix C Metric Definitions
For a ground truth depth map $z$ and a predicted $\hat{z}$, AbsRel is defined as the average of $|z_{i}-\hat{z_{i}}|/z_{i}$. Following Wang et al. (2025a), we define AbsRelp on point maps $p\in\mathbb{R}^{H\times W\times 3}$ by the average of $\left\|p_{i}-\hat{p}_{i}\right\|/\|p_{i}\|$.
@@
$\text{Log RMSE SI}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(\log z_{i}-\log\hat{z_{i}}+\alpha)^{2}}\quad\text{ where }\alpha=\frac{1}{n}\sum_{i}(\log\hat{z}_{i}-\log z_{i}).$
This follows the definitions in Ke et al. (2024); Eigen et al. (2014).
Finally, WKDR follows the definition given in Yin et al. (2021) based on relative depth relationships between pairs of pixels. Boundary F1 score follows the definition in Bochkovskii et al. (2025) to compute an F1 score over the edges in the predicted and ground truth depth maps.
The affine and scale alignment procedure used throughout is identical to MoGe (Wang et al. (2025a)). Notably, affine alignment for the point cloud metric AbsRelp is not the same as affine alignment of the depth. “Lst. Sq.” denotes least squares alignment on depth, as performed by Marigold (Ke et al. (2024)).
Appendix D Relative Normal
D.1 Sampling Procedure
@@
When computing the relative normal metric, we wish to ensure that our deterministic sampling technique accurately approximates the true mean of the distribution. In Fig. 9 we display the difference between the metric computed with the deterministic algorithm and the metric computed with $10^{8}$ random samples. We perform the error computations using predictions from Depth Pro (Bochkovskii et al. (2025)), UniDepthV2 (Piccinelli et al. (2025a)), MoGeV1 (Wang et al. (2025a)), Metric3DV2 (Hu et al. (2024)), and MoGeV2 (Wang et al. (2025b)) on 100 images from the iBims dataset and 100 images from the Virtual KITTI dataset Koch et al. (2019); Cabon et al. (2020). The maximum error $5.84\times 10^{-4}$ and average error of $1.08\times 10^{-4}$ is sufficiently small to justify using $10^{6}$ samples. We also analyze the $z$-score of the deterministic computation compared to random sampling. This suggests that using the Sobol sequence does not introduce unexpected irregularities.
Appendix E SAWA-H Comparison Examples
@@ Appendix F Dataset Gallery
Appendix G Performance of MDE Methods under SAWA-H

### v3

@@ 5 Sensitivity Aligned Composition
Our sensitivity analysis of existing metrics reveals that no existing metrics align well with human judgment. How do we create a metric that does align well, if one is desired for some reason?
@@
Similar to WKDR ([zoran2015ordinalrelationship]) and boundary F1 score ([bochkovskiydepthpro]), we consider the relation between patches (sets of pixels) within the depth map. Given patches $p$ and $q$ we calculate the normal of the surface described by each patch and the angle between these normals $\angle(\hat{n}_{p},\hat{n}_{q}),\angle(n_{p},n_{q})$ for both ground truth and prediction. We take the average difference over many such patches,
@@
Our selection of patches is motivated by the following procedure: first uniformly select a pixel $I$ within the image, then select a pixel $J$ uniformly within a neighborhood of $I$, finally compute the error. To approximate this expectation over a multivariate uniform distribution, we use the first $m$ elements of the Sobol sequence. This provides a deterministic algorithm for computing the metric. In Appendix 12 we show that 1 million samples is sufficient when computing this metric on standard RGB-D datasets. We take the average over multiple scales by downsampling the image by a factor $k$ and computing the relative normals. The following results use a neighborhood radius of 32 and scale factors $[1,2,4,8]$. We compute normals using the cross product of vectors from the top/bottom and left/right pixels adjacent to a central pixel. For datasets with noisy ground truth, it may be preferable to compute normals with a plane of best fit.
@@
Note that we do not claim that SAWA-H is a better metric, or a substitute for any existing metric. We only claim that SAWA-H aligns better with human judgment, which may or may not be desirable depending on the use case. Also, our alignment approach is not restricted to human preferences. One can create composite metrics to align with arbitrary preferences based on the desired sensitivity to perturbations.
@@
Example of SAWA-H on Real Predictions In some cases, the SAWA-H (with RelNormal) metric provides different rankings among depth predictions than the AbsRel metric. In Fig. 6 we visualize one such difference for the predictions of Depth Pro and Metric3DV2 on an Infinigen scene using error maps. This is mostly straightforward, though to compute the error map for the boundary f1 score we divide the error equally among all edge pixels where the ground truth and predicted depth map disagree. Though Metric3D achieves a smaller AbsRel error, its SAWA-H error is substantially larger. This is due to the fact that the Depth Pro model made a mistake judging the scale of the iceberg and orientation of the camera relative to ocean. Hence, the iceberg and ocean cannot both be properly aligned, so the global reconstruction is inaccurate and the AbsRel error is high. As seen in the projected contour plot, Depth Pro better approximates the local geometry and has a lower SAWA-H error score. We present more examples with predictions on Infinigen and iBims [ibims] in Appendix  13. It is important to note that this example does not show that SAWA-H is a better metric than AbsRel because that depends on the application; it simply illustrates that their different behaviors.
Performance of MDE Methods under SAWA-H We use SAWA-H (with RelNormal) to evaluate 10 state-of-the-art monocular depth estimation methods: DepthAnything[yang2024depthanything], DepthAnythingV2[yang2024depthanythingv2], GeoWizard[fu2024geowizard], Marigold[ke2024marigold], MiDaS[birkl2023midas], DepthPro[bochkovskiydepthpro], Metric3DV2[hu2024metric3d], MoGe-2[wang2025moge2], UniDepthV2[piccinelli2025unidepthv2], and ZoeDepth[bhat2023zoedepth]. Evaluation is conducted on 9 commonly used benchmarks: DDAD[guizilini2020DDAD], DIODE[vasiljevic2019diode], ETH3D[schops2019bad_eth3d], HAMMER[jung2023hammer], iBims-1[ibims], KITTI[uhrig2017sparsity_kitti], NYUv2[silberman2012indoor_nyuv2], Sintel[butler2012naturalistic_sintel] and Spring[mehl2023spring].
@@
Fig. 5 presents metric values of each method averaged over 9 datasets. Performance of methods on each dataset are reported in Appendix 15. Note that Boundary F1 is one component of SAWA-H but some datasets do not provide valid depth value on pixels near boundaries, preventing SAWA-H from properly penalizing blurry boundaries on those datasets. We additionally report performance of methods averaged on datasets with accurate depth across boundaries in Appendix 15.
@@ 7 Acknowledgments
[…]
Appendix
8 Perturbation Algorithm Of Controllable Intensity
@@
9 More on Collecting Data of Human Judgment
@@ 10 Metric Definitions
For a ground truth depth map $z$ and a predicted $\hat{z}$, AbsRel is defined as the average of $|z_{i}-\hat{z_{i}}|/z_{i}$. Following [wang2025moge], we define AbsRelp on point maps $p\in\mathbb{R}^{H\times W\times 3}$ by the average of $\left\|p_{i}-\hat{p}_{i}\right\|/\|p_{i}\|$.
@@
$\text{Log RMSE SI}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(\log z_{i}-\log\hat{z_{i}}+\alpha)^{2}}$
where $\alpha=\frac{1}{n}\sum_{i}(\log\hat{z}_{i}-\log z_{i})$.
This follows the definitions in [ke2024marigold, eigen2014depth].
Finally, WKDR follows the definition given in [yin2021learning] based on relative depth relationships between pairs of pixels. Boundary F1 score follows the definition in [bochkovskiydepthpro] to compute an F1 score over the edges in the predicted and ground truth depth maps.
The affine and scale alignment procedure used throughout is identical to MoGe ([wang2025moge]). Notably, affine alignment for the point cloud metric AbsRelp is not the same as affine alignment of the depth. “Lst. Sq.” denotes least squares alignment on depth, as performed by Marigold ([ke2024marigold]).
11 Causes of Different Behaviors Between Human Judgment and Existing Metrics
Fig. 7 shows instances of relative scale, surface orientation, camera intrinsics, and curvature perturbations. On each scene, instance of curvature perturbation has much lower (better) AbsRel score, but its shape is very different from ground truth.
Fig. 8 shows instances of affine transform perturbations. The shape of these instances is distorted a lot, but they achieve a perfect score when the respective type of affine alignment is applied during evaluation.
12 Relative Normal
12.1 Sampling Procedure
@@
When computing the relative normal metric, we wish to ensure that our deterministic sampling technique accurately approximates the true mean of the distribution. In Fig. 9 we display the difference between the metric computed with the deterministic algorithm and the metric computed with $10^{8}$ random samples. We perform the error computations using predictions from Depth Pro ([bochkovskiydepthpro]), UniDepthV2 ([piccinelli2025unidepthv2universalmonocularmetric]), MoGeV1 ([wang2025moge]), Metric3DV2 ([hu2024metric3d]), and MoGeV2 ([wang2025moge2]) on 100 images from the iBims dataset and 100 images from the Virtual KITTI dataset [ibims, vkitti]. The maximum error $5.84\times 10^{-4}$ and average error of $1.08\times 10^{-4}$ is sufficiently small to justify using $10^{6}$ samples. We also analyze the $z$-score of the deterministic computation compared to random sampling. This suggests that using the Sobol sequence does not introduce unexpected irregularities.
13 SAWA-H Comparison Examples
@@ 14 Dataset Gallery
15 Performance of MDE Methods under SAWA-H